OpenStudy (anonymous):

How to find the indicated derivative of 6t^3+3/t^4 when all my answer choices are two fractions???

OpenStudy (anonymous):

\[f(t) = 6t^{3} + \frac{3}{t^{4}} = 6t^{3} + {3}{t^{-4}}\] Do chain rule to get f'(t) \[f(x) = a*g(x)^{n}\] \[f'(x) = (n-1)*a*g(x)^{n-1}*g'(x)\]

OpenStudy (anonymous):

f'(x)= 12t-12t^-5

OpenStudy (anonymous):

the derivative of 6t^3 is 3*6t^(3-1) = 18t^2

OpenStudy (anonymous):

so i got 18/t^2 - 12/t^5 ?

OpenStudy (anonymous):

18t^2 - 12/t^5 that '/' doesn't need to be in there

OpenStudy (turingtest):

your problem is\[6t^3+\frac3{t^4}\]or\[\frac{6t^3+3}{t^4}\]?

OpenStudy (anonymous):

the first one but all my answer choices are fractional like the one i posted in reply to 2Le

OpenStudy (anonymous):

they are simplifying it weird..

OpenStudy (turingtest):

very weird

OpenStudy (anonymous):

is the answer \[\frac{18t^7 -12}{t^5}\]

OpenStudy (turingtest):

the derivative is\[18t^2-\frac{12}{t^5}\]which you could write with a common denominator if you want to be stupid about it :P

OpenStudy (turingtest):

\[18t^2-\frac{12}{t^5}=\frac{18t^7-12}{t^5}\]is that a choice?

OpenStudy (anonymous):

no, the choices look like: -6/t - 12/t^5 and 4 others in similar format and signs are all the same but some powers vary and some numerators

OpenStudy (turingtest):

well the -12/t^5 part is fine, but something is screwy with the first term

OpenStudy (anonymous):

oh sorry, it actually looks like the second one -_- haha. (your first post)

OpenStudy (turingtest):

sweet :)

OpenStudy (anonymous):

everything under t^4...

OpenStudy (anonymous):

those solution manuals and answers in the back are usually written by some sort of computer. Check your work and reformatting it is usually a second part of the problem.

OpenStudy (anonymous):

it is a computerized old quiz i am reviewing for a test and its the only one im stumped about